# Explainable Molecular Property Prediction: Aligning Chemical Concepts with Predictions via Language Models

### v2

@@ 4.1 Experimental Setup
Datasets. We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction. Six mutagenicity datasets are Mutag [37], Mutagen [38], PTC-FM [39], PTC-FR [39], PTC-MM [39], and PTC-MR [39]. For hepatotoxicity [39], the Liver dataset [40] is used. We used these datasets as the ground truth substructures in these datasets are known. For detailed information on these datasets, please refer to Appendix 2.
Evaluation Metrics. We use three metrics for evaluating the performance of the proposed Lamole.
  1. 1.
    Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$.
  2. 2.
    Explanation Accuracy: We follow the experimental settings in GNNExplainer [12], which formulates the explanation problem as a binary classification of edges. We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation. We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones.
  3. 3.
    Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
@@
GCN [41] 84.6 78.9 54.8 63.0 57.8 53.3 41.1
DGCNN [42] 85.8 74.8 57.3 63.5 61.0 58.6 44.6
edGNN [43] 86.9 75.2 59.8 65.7 64.4 56.3 44.5
GIN [44] 87.5 82.3 62.1 66.2 65.1 64.0 44.9
RW-GNN [45] 87.2 80.3 61.9 64.0 62.4 57.0 43.2
DropGNN [46] 89.4 80.7 62.0 66.0 63.7 64.2 45.0
IEGN [47] 84.6 80.1 60.8 59.8 61.1 59.5 45.3
ChemBERTa [10] 86.8 78.0 60.0 65.7 60.4 58.7 45.7
@@
  • and denote the best result and second result, respectively.
Baselines. We combine Lamole into three BERT family models, such as DistilBert [48], DeBerta [49], and Bert to evaluate the performance of the proposed Lamole. For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa [10] and several GNNs including GCN [41], DGCNN [42], edGNN [43], GIN [44], RW-GNN [45], DropGNN [46], and IEGN [47].
For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad [50], CAM [51], GradInput [52], and GradCAM [13], 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX [11], GNNExplainer [12], PGExplainer [21], and Gem [20].
Implementation Details. Following ChemBERTa [10], our used BERT family models are pre-trained on 100,000 molecules from the ZINC dataset [53]. 15% tokens in each input string are randomly for masked language learning. We finetuned the models on Mutag, Mutagen, PTC-FM, PTC-FR, PTC-MM, PTC-MR, and Liver. For each dataset, the ratio of samples with annotation masks over the data size is 10%. The details of experimental settings can be found in Appendix 3.
4.2 Results
@@
SmoothGrad+GCN [50] 68.0 66.1 66.8 62.2 64.5 61.9 66.0
CAM+GCN [51] 66.8 65.4 66.9 62.7 61.4 65.5 67.5
GradInput+GCN [52] 69.1 65.1 68.5 64.3 62.3 62.0 67.2
GradCAM+GCN [13] 68.3 65.7 70.4 65.4 65.8 66.6 68.4
SmoothGrad+Bert [50] 71.3 71.3 72.8 70.6 65.2 68.9 74.4
CAM+Bert [51] 70.1 69.8 66.3 69.4 63.9 66.2 72.8
GradInput+Bert [52] 73.1 71.1 72.1 66.6 64.0 68.1 75.0
GradCAM+Bert [13] 73.5 71.5 76.3 69.0 69.1 71.2 75.8
OrphicX [11] 76.4 70.4 71.1 68.4 69.0 69.1 73.0
GNNExplainer [12] 69.5 63.0 68.6 67.3 66.0 66.3 71.4
PGExplainer [21] 65.7 58.0 69.3 66.3 64.7 65.1 71.0
Gem [20] 72.8 65.5 69.8 66.8 68.2 68.3 72,5
@@
Figure 5: The explanation accuracy of Lamole with different annotation rates on the Mutag, PTC-FM, and PTC-MR datasets.
(a) Mutag
(b) PTC-FM
(c) PTC-MR
@@
Molecular Property Prediction. Table 1 shows the classification accuracy of compared algorithms. Our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert slightly outperform most compared algorithms on the Mutag, PTC-FM, PTC-FR, PTC-MR, and Liver datasets. Though the LMs, including Lamole and ChemBERTa, demonstrate good performance for molecular property prediction, they are inferior to GNN-based methods on the Mutagen and PTC-MM datasets. In addition, Lamole models show superior performance over ChemBERTa. This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the structure-property relationships.
Explanation Performance. Table 2 presents the explanation accuracy of the compared explainability techniques. It is obvious that the LM-based methods using the Group SELFIES strings demonstrate significantly higher explanation accuracy over explainable GNNs. Specifically, Lamole improves the explanation accuracy by $1.4\%\sim 14.8\%$ compared to the baseline methods. This is because the molecular graph-based explainable GNNs can only highlight individual atoms and bonds rather than chemically meaningful substructures like functional groups. Instead, Group SELFIES explicitly encode the chemically meaningful substructures as tokens, and using Group SELFIES can help explainable methods to generate chemically meaningful explanations by highlighting the corresponding tokens.
@@
We also selected some representative molecules for explanation visualization. These explanations are shown in Figs. 1, 4, and 12$\sim$15 (see Appendix), respectively. The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments. As shown in Fig. 4, our proposed Lamole assigns the highest two importance scores to the two ground truth function groups, i.e., benzene-1 and nitro group. In contrast, OrphicX cannot provide chemically meaningful explanations, and CradCAM+Bert cannot accurately identify the functional group interactions. Compared to baseline methods, Lamole provides chemically meaningful explanations. Particularly, the interaction among the functional groups is successfully captured. The visualization demonstrates Lamole’s superior interpretation in uncovering the structure-property relationships.
We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%), as presented in Figs. 5 and 10 (see Appendix). Clearly, using more molecules with ground truth annotations can continuously improve explanation accuracy. In addition, compared to the results without using the marginal loss $\mathcal{L}_{M}$, using only 10% molecules with ground truth annotations can significantly improve explanation accuracy by up to 5%, confirming the effectiveness of using annotations. Though increasing the annotated ground truth rate can only bring a limited increase in the explanation accuracy, there is a trade-off between the achieved explanation accuracy and the additional annotation costs.
@@
In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$. The statistical results of $\mathrm{EP}$ are presented in Figs. 6 and 11 (see Appendix). From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property. Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$. The statistical results of $\mathrm{EP}$ in Figs. 6 and 11 (see Appendix) indicate Lamole can provide explanations with higher confidence in matching the ground truth substructures.
Due to the page limitation, more analysis regarding the ablation studies on attention weights and the marginal loss, the impact of annotation rates, the analysis regarding the explanation plausibility, and more visualization results can be found in the Appendix.
@@ 5 Conclusions
This work developed a new framework for explainable molecular property prediction based on language models, called Lamole. Lamole uses Group SELFIES as input for perceiving the chemically meaningful semantics. By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations that can capture the functional group interactions. Furthermore, one marginal loss is designed to calibrate the explanations by aligning them with the chemists’ annotation. With the above designs, Lamole can provide chemically meaningful explanations while faithfully uncovering the structure-property relationships. Lamole ’s effectiveness has been demonstrated through theoretical analysis and extensive experimental validation.

### v3

@@ 4.1 Experimental Setup
Datasets. We use six datasets on two types of tasks, i.e., hepatotoxicity and mutagenicity, to evaluate the algorithmic performance for the explainable molecular property prediction. Six mutagenicity datasets are Mutag (Debnath et al., 1991), Mutagen (Morris et al., 2020), PTC-FM (Toivonen et al., 2003), PTC-FR (Toivonen et al., 2003), PTC-MM (Toivonen et al., 2003), and PTC-MR (Toivonen et al., 2003). For hepatotoxicity (Toivonen et al., 2003), the Liver dataset (Liu et al., 2015) is used. We used these datasets as the ground truth substructures in these datasets are known. For detailed information on these datasets and the ground truth substructures, please refer to Appendix.
Evaluation Metrics. We use three metrics for evaluating the performance of the proposed Lamole. 1) Classification Accuracy: We evaluate the model’s predictions by $\sum_{i=1}^{I}\mathbb{I}(y^{(i)}=\hat{y}^{(i)})/I$. 2) Explanation Accuracy: We follow the experimental settings in GNNExplainer (Ying et al., 2019), which formulates the explanation problem as a binary classification of edges. We treat edges inside ground-truth substructure as positive edges and negative otherwise, and AUC is adopted as the metric for quantitative evaluation. We only consider the mutagenic/hepatotoxic molecules because no explicit substructures exist in nonmutagenic/nonhepatotoxic ones. 3) Explanations’ Plausibility: We use the defined explanations’ plausibility $\mathrm{EP}$ to measure how confident the explanation aligns with the ground truth.
Baselines. We combine Lamole into three BERT family models, such as DistilBert (Sanh, 2019), DeBerta (He et al., 2020), and Bert to evaluate the performance of the proposed Lamole. For evaluating classification accuracy, we compare our Lamole with one SMILES string-based LM, ChemBERTa (Chithrananda et al., 2020) and several GNNs including GCN (Kipf & Welling, 2016), DGCNN (Zhang et al., 2018), edGNN (Jaume et al., 2019), GIN (Xu et al., 2018), RW-GNN (Nikolentzos & Vazirgiannis, 2020), DropGNN (Papp et al., 2021), and IEGN (Maron et al., 2018).
For evaluating explanation accuracy, Lamole is compared with three types of alternative methods: 1) GCN with feature-based explainability techniques, including SmoothGrad (Smilkov et al., 2017), GradInput (Shrikumar et al., 2017), and GradCAM (Selvaraju et al., 2017), 2) Bert with the above feature-based explainability techniques, where Group SELFIES is used as input for a fair comparison, and 3) explainable GNNs including OrphicX (Lin et al., 2022), GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), and Gem (Lin et al., 2021). The details of experimental settings can be found in Appendix.
4.2 Results
Prediction Performance. Table 1 shows the classification accuracy of compared algorithms. As we can see, our proposed Lamole+DistilBert, Lamole+DeBerta, and Lamole+Bert not only can provide explainability but also can achieve comparable prediction accuracy as compared to existing predictive methods. In addition, Lamole models show superior performance over ChemBERTa. This suggests that using molecular representations with more chemical semantics, like Group SELFIES, can help LMs better learn the chemical semantics and structure-property relationships.
@@ 4.2 Results
GCN (Kipf & Welling, 2016) 84.6 78.9 54.8 63.0 57.8 53.3 41.1
DGCNN (Zhang et al., 2018) 85.8 74.8 57.3 63.5 61.0 58.6 44.6
edGNN (Jaume et al., 2019) 86.9 75.2 59.8 65.7 64.4 56.3 44.5
GIN (Xu et al., 2018) 87.5 82.3 62.1 66.2 65.1 64.0 44.9
RW-GNN (Nikolentzos & Vazirgiannis, 2020) 87.2 80.3 61.9 64.0 62.4 57.0 43.2
DropGNN (Papp et al., 2021) 89.4 80.7 62.0 66.0 63.7 64.2 45.0
IEGN (Maron et al., 2018) 84.6 80.1 60.8 59.8 61.1 59.5 45.3
ChemBERTa (Chithrananda et al., 2020) 86.8 78.0 60.0 65.7 60.4 58.7 45.7
@@
  • and denote the best and the second-best results, respectively.
Explanation Performance. Table 2 presents the explanation accuracy of the compared explainability techniques. It should be noted that the ground truth annotations used in our work provide additional supervisory signals. Therefore, we also align the generated explanations of these baselines with the annotations for fair comparison, and the ground truth annotation rate is 10%. Lamole improves the explanation accuracy by $1.4\%\sim 14.3\%$ compared to the baseline methods. We also investigated the explanation accuracy of our Lamole under different ground truth annotation rates (10%, 20%, 50%, and 100%). The impact of annotation rates can be found in Appendix. In addition, we discuss the rationale of using labeled annotations (see Appendix).
@@
GradInput+GCN (Shrikumar et al., 2017) 70.3 67.9 69.7 66.4 64.6 65.0 73.0
GradCAM+GCN (Selvaraju et al., 2017) 69.8 67.0 71.0 67.9 66.2 67.3 69.4
SmoothGrad+GCN (Smilkov et al., 2017) 69.2 66.8 67.5 62.6 64.9 63.1 66.4
GradInput+Bert (Shrikumar et al., 2017) 75.1 72.6 73.0 68.9 65.6 69.6 75.6
GradCAM+Bert (Selvaraju et al., 2017) 75.3 72.4 77.5 70.0 70.2 73.0 76.0
SmoothGrad+Bert (Smilkov et al., 2017) 73.4 72.8 73.7 71.0 67.0 69.9 75.1
GNNExplainer (Ying et al., 2019) 70.6 64.2 68.9 67.9 66.8 67.1 72.1
PGExplainer (Luo et al., 2020) 66.5 58.7 70.3 68.0 65.9 67.0 71.5
Gem (Lin et al., 2021) 73.7 66.0 71.3 69.0 68.9 69.2 73.6
OrphicX (Lin et al., 2022) 78.0 71.5 74.6 70.4 70.9 71.4 74.0
@@
  • and denote the best and the second-best results, respectively.
@@
We selected some representative molecules for explanation visualization. These explanations are shown in Figs. 1 and 4, respectively. The right panel of those figures is the importance scores obtained by Lamole, where "other" in the figures is the average importance score of other unlisted functional groups/fragments. Compared to baseline methods, Lamole provides chemically meaningful explanations. Particularly, the interaction among the functional groups is successfully captured. More visualization results can be found in Appendix.
@@
In addition, we evaluated the performance of compared algorithms by using the proposed explanation plausibility metric $\mathrm{EP}$. The statistical results of $\mathrm{EP}$ are presented in Figs. 5 and LABEL:epfr. From the figures, we can observe that the $\mathrm{EP}$ values of the comparison algorithm are slightly lower, which means that the algorithms cannot confidently reflect the relationships between structure and property. Compared to the comparison algorithm, the $\mathrm{EP}$ values of Lamole have increased by $2\%\sim 9\%$. More analysis regarding the explanation plausibility can be found in Appendix.
4.3 Ablation Studies
We conducted ablation studies for each component in our Lamole. Specifically, we removed the attention weights in the explanations, removed the marginal loss, and only used attention weights as explanations. The corresponding ablation algorithms are named Lamole ($-\alpha$), Lamole ($-\mathcal{L}_{M}$), and Lamole (Att), respectively. The results are shown in Fig. 6. The explanation accuracy decreases by 1.4$\sim$2.3when removing the attention weights. Removing the marginal loss can decrease the explanation accuracy by 1.0$\sim$5.0%. Regarding the results of using only attention weights, The explanation accuracy decreases by 1.4$\sim$6.3%. The above results confirm the effectiveness of using the marginal loss, attention weights, and gradients.
Figure 6: The explanation accuracy of Lamole+Bert, Lamole+Bert ($-\alpha$), Lamole+Bert ($-\mathcal{L}_{M}$), and Lamole (Att).
From the perspective of model training, the marginal loss enables the model to be trained under the causal signals. To verify, we compared the classification performance with and without the marginal loss, as shown in Table 3. Without the marginal loss, the classification accuracy degrades by 0.7$\sim$3.7%. The above results indicate that marginal loss could help identify the causal features, thereby improving classification accuracy.
Table 3: The classification performance with and without the marginal loss (%)
Methods Mutag Mutagen PTC-FM PTC-FR PTC-MM PTC-MR Liver
Lamole-$\mathcal{L}_{M}$ 84.9 73.8 58.6 69.3 59.7 62.3 44.6
Lamole 88.2 74.5 62.4 70.0 61.2 66.0 47.5
To investigate the attention weights, the attention weights of two molecules are depicted in Fig. 7. From Fig. 7 middle panel, it can be found the correlation among the ground truth substructures is higher than others, showcasing the rationality of using attention weights to capture the functional group interactions. When we aggregate the attention weights for each token, as shown in Fig. 7 right panel, the top two attention weights of the molecule (ID:155) match the two ground truth substructures. However, the top two attention weights of the molecule (ID:156) do not match the two ground truth substructures. The above results indicate that attention weights can capture the interactions and also confirm that "attention is not explanation" (Jain & Wallace, 2019; Serrano & Smith, 2019; Abnar & Zuidema, 2020). Due to the page limitation, the limitation of the proposed work is provided in Appendix.
Figure 7: Attention weights visualization. The ground truth substructures are highlighted in red.
@@ 5 Conclusions
This work proposed Lamole for explainable molecular property prediction based on language models. Lamole uses Group SELFIES as input for chemically meaningful semantics. By disentangling the information flows of Transformer-based LMs, Lamole integrates attention weights into gradients to generate explanations to quantify each chemically meaningful substructure’s impact on the model’s output. Furthermore, one marginal loss is designed to calibrate the explanations to be more faithful by aligning them with the chemists’ annotation. Lamole ’s effectiveness has been demonstrated through theoretical analysis and extensive experimental validation.